# SepLLM: Accelerate Large Language Models by Compressing One Segment into One Separator

### v5

@@ 1 Introduction
The core challenge underlying these efficiency issues is the self-attention module, which exhibits quadratic complexity with respect to the number of input tokens. Research on efficient Transformers in LLMs primarily follows two major directions. The first approach focuses on linear attention (Katharopoulos et al., 2020; Schlag et al., 2021), replacing the vanilla self-attention module with alternatives that achieve linear complexity. However, these modifications make the architecture significantly different from traditional self-attention, preventing direct utilization of powerful pre-trained Transformer models. The second approach emphasizes KV cache optimization (Xiao et al., 2024a; Zhu et al., 2024; Xiao et al., 2024b; Li et al., 2024b), aiming to eliminate redundant KV cache to accommodate longer input contexts. For example, Xiao et al. (2024a) introduced an adaptive mechanism that selectively retains essential tokens and their KV based on cumulative attention scores. Similarly, Zhu et al. (2024) proposed a token selection strategy with controlled sparsity, achieving near-lossless acceleration. While promising, these training-free methods adapt poorly to the training stage, resulting in discrepancies between training and inference performance. StreamingLLM (Xiao et al., 2024b) represents a notable attempt to address these limitations by preserving attention sinks and local tokens to reduce computation and memory overhead. However, it omits many intermediate tokens, resulting in performance degradation compared to standard Transformers.
@@
  • We have made our implementation publicly available at sepllm.github.io. Our codebase supports efficient multi-node distributed training with accelerated attention module Sep-Attention and also supports numerous existing Fusion Operators to accelerate the training process, such as fused rope (Su et al., 2023), fused layer norm, etc.
@@ 2 Related Work
KV Cache Compression. Recent research has focused on overcoming LLMs’ limitations in processing extensive contextual inputs. FastGen (Ge et al., 2024) proposes an adaptive KV cache management method, optimizing memory usage by customizing retention strategies for different attention heads. SnapKV (Li et al., 2024b) enhances efficiency through KV cache compression, utilizing attention scores to select and cluster significant positions. H2(Zhang et al., 2024b) implements a dynamic token retention policy, balancing recent and historically important information to optimize memory use. StreamingLLM (Xiao et al., 2024b) expands LLMs’ capabilities to handle infinite sequence lengths without fine-tuning, by reserving attention sinks and local tokens. QuickLLaMA (Li et al., 2024a) proposes to evict the query-aware KV cache for inference acceleration. PyramidInfer (Yang et al., 2024) and PyramidKV (Zhang et al., 2024a) modify the KV cache capacity across different layers, prioritizing larger allocations in the lower layers while reducing those in the upper layers. However, most works in this category cannot be applied into training phase.
@@ 3.1 Fundamental Design
Separator Tokens. From Figure 2, we can observe that within a given input context, seemingly “meaningless” separator tokens (such as commas, periods, exclamation marks, semicolons, etc.) that segment sequences receive higher attention scores compared to semantically meaningful tokens (such as nouns or verbs). Therefore, we hypothesize that these separators may compress the information of the text segments naturally segmented by them, such that when the Transformer generates new tokens, it only needs to reference the information contained in these separators to extract the information pertaining to those text segments. Hence, in a training-free scenario, we employed this strategy and achieved similar results to the original model based on full attention across many tasks. Furthermore, to reinforce the effect of using separators to compress information within their respective segments, we employed training-from-scratch and post-training approaches to compel the model during training to restrict the current token from accessing all information from distant preceding text, i.e., in each segment, only the separator representing its segment is visible to the current token (with other tokens being masked, see Figure 3). After training in this manner, the information within segments is forced to be condensed into the separators, leading the Transformer’s probability distribution for predicting the next word closely resembling that of the original Transformer with full attention. See more in Appendices F and G.
Neighboring Tokens. Language tasks usually exhibit strong local dependencies and interactions, since adjacent tokens often form coherent phrases or have dependencies that are required to be captured. Neighboring tokens usually help form locally smooth and coherent contexts, allowing the model to generate sentences that are reasonable within the immediate context. The neighboring tokens, also referred to as local attention or sliding-window attention, are considered in various efficient Transformers (Xiao et al., 2024b; Zhang et al., 2024b) and we have also adopted this approach, with the number of preceding tokens closest to the current token denoted as “n”.
@@ 3.2 Overall Pipeline
We split the overall pipeline of our proposed SepLLM into training/pre-filling stage and generating stage. We also provide a theoretical analysis about the Universal Approximation of SepLLM in Appendices H and I.
@@ 4.1 Experimental Settings
Table 1: Evaluation results and average runtime KV cache usage for training-free experiments on GSM8K-CoT 8-shots and MMLU 5-shots. For SepLLM and StreamingLLM, three initial tokens’ KV are kept for this experiment. r.KV(%) here represents the ratio of KV usage at runtime for the respective method compared to Vanilla. See more results in Appendices G and Table 14.
@@ 4.2 Training-free
Method ARC-c ARC-e LBD-ppl LBD-acc LogiQA PIQA SciQ Attn(%) r.KV(%)
@@ Results.
The experimental results for training-free are shown in Table 1. “Vanilla” represents the original Llama-3 model with full attention, while “StrmLLM” represents StreamingLLM (Xiao et al., 2024b). n means the number of KV for Neighboring Tokens we retain. For SepLLM, all the KV for Separator Tokens are kept and for the setting SepLLM (n=256), we find that SepLLM exhibits comparable performance in both multi-step mathematical CoT task and multidisciplinary knowledge reasoning tasks, when compared to the full-attention Llama-3. SepLLM achieves this using only 47.36% of the KV utilized by the original Llama-3 for reasoning, indicating SepLLM’s capability of modeling both the contexts requiring multi-step logical analysis and those involving multi-domain knowledge reasoning while retaining only 50% original KV.
@@ 4.6 Ablation Study
We also perform the following experiments to validate the effectiveness of Initial Tokens and PE’s shifting for both StreamingLLM and SepLLM. The experimental results are shown in Table 8 and the discussions are as follows.
@@ 4.7 Generalization and Information Retrieval
To verify the generalization of SepLLM, we adapt SepLLM to models of different architectures and scales. The results in Appendix D can validate the generalization of our proposed SepLLM. Specifically, we adapt SepLLM to different backbones including Pythia-6.9B, Pythia-12B (Biderman et al., 2023), Llama-3-8B-Base/Instruct (Dubey et al., 2024) and Falcon-40B (Almazrouei et al., 2023). Moreover, we also conduct the Needle In A Haystack experiment, which further demonstrates the compression capability of separator tokens for segment information. As illustrated in Appendix E, SepLLM can retrieve the needle in most scenarios. In comparison, StreamingLLM (Xiao et al., 2024b) cannot complete this task. Appendices F and G discuss the effect of separators. Appendices H and I provide a theoretical analysis on the Universal Approximation of SepLLM.
@@ Appendix D The Performance of Different Models
Appendix E Needle In A Haystack
@@
Figure 8: Needle In A Haystack test results for streamingLLM (n=64) based on Pythia-160M-deduped.
Figure 9: Needle In A Haystack test results for our SepLLM(n=64, H/T) based on Pythia-160M-deduped.
Figure 10: Needle In A Haystack test results for our SepLLM(n=2048; first/last 2 layers (4 layers in total): full attention) based on Llama-3-8B-instruct. 4 initial tokens are kept.
Figure 11: Needle In A Haystack test results for our SepLLM(n=2048; first/last 2 layers (4 layers in total): full attention) based on Llama-3-8B-instruct. 32 initial tokens are kept.
@@
Appendix F Discussions on Separators
@@
Appendix G Fixed-interval Variant
To further verify the importance of separators, we propose another variant, i.e., when calculating sparse attention, instead of focusing only on Separator Tokens between the Initial Tokens and Neighboring Tokens, we attend to one token at fixed intervals (e.g., every 8 tokens or 16 tokens), while the other tokens are masked, namely FixLLM. We evaluated the mathematical reasoning ability and knowledge-based reasoning ability of the two variants using the GSM8K-CoT (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021)) benchmarks under a training-free setting, based on the Llama3-8B-Instruct backbone. See results in Table 14.
@@ Appendix G Fixed-interval Variant
Table 14: Evaluation results and average runtime KV cache usage for training-free experiments on GSM8K-CoT 8-shots and MMLU 5-shots. For SepLLM and FixLLM, three initial tokens’ KV are kept. $\Delta$ denotes the interval size for FixLLM and n is the number of retained neighboring tokens’ KV. r.KV(%) here represents the ratio of KV usage at runtime for the respective method compared to Vanilla.
Appendix H Universal Approximation
@@
Appendix I Proof for Theorem H.1
Lemma I.1 (Lemma 5 in Yun et al. (2020)).
@@
Definition I.3 (Contextual Mapping).
@@
Lemma I.5 (Lemma 8 in Yun et al. (2020)).
For the contextual mapping $g_{c}$ in Lemma I.4, there exist $\frac{n}{\delta^{dn}}$-layer modified SepLLM (composed of feed-forward layers and identity attention layers), denoted as $g_{v}\in\bar{\mathcal{T}}^{2,1,1}_{\text{Sep}}:\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times n}$, such that

### v6

@@ 1 Introduction
The core challenge underlying these efficiency issues is the self-attention module, which exhibits quadratic complexity with respect to the number of input tokens. Research on efficient Transformers in LLMs primarily follows two major directions. The first approach focuses on linear attention (Katharopoulos et al., 2020; Schlag et al., 2021), replacing the vanilla self-attention module with alternatives that achieve linear complexity. However, these modifications make the architecture significantly different from traditional self-attention, preventing direct utilization of powerful pre-trained Transformer models. The second approach emphasizes KV cache optimization (Xiao et al., 2024a; Zhu et al., 2024; Xiao et al., 2024b; Li et al., 2024), aiming to eliminate redundant KV cache to accommodate longer input contexts. For example, Xiao et al. (2024a) introduced an adaptive mechanism that selectively retains essential tokens and their KV based on cumulative attention scores. Similarly, Zhu et al. (2024) proposed a token selection strategy with controlled sparsity, achieving near-lossless acceleration. While promising, these training-free methods adapt poorly to the training stage, resulting in discrepancies between training and inference performance. StreamingLLM (Xiao et al., 2024b) represents a notable attempt to address these limitations by preserving attention sinks and local tokens to reduce computation and memory overhead. However, it omits many intermediate tokens, resulting in performance degradation compared to standard Transformers.
@@
  • We have made our implementation publicly available at sepllm.github.io. Our codebase supports efficient multi-node distributed training with accelerated attention module Sep-Attention and also supports numerous existing Fusion Operators to accelerate the training process, such as fused rope (Su et al., 2024), fused layer norm, etc.
@@ 2 Related Work
KV Cache Compression. Recent research has focused on overcoming LLMs’ limitations in processing extensive contextual inputs. FastGen (Ge et al., 2024) proposes an adaptive KV cache management method, optimizing memory usage by customizing retention strategies for different attention heads. SnapKV (Li et al., 2024) enhances efficiency through KV cache compression, utilizing attention scores to select and cluster significant positions. H2(Zhang et al., 2023) implements a dynamic token retention policy, balancing recent and historically important information to optimize memory use. StreamingLLM (Xiao et al., 2024b) expands LLMs’ capabilities to handle infinite sequence lengths without fine-tuning, by reserving attention sinks and local tokens. QuickLLaMA (Li et al., 2025) proposes to evict the query-aware KV cache for inference acceleration. PyramidInfer (Yang et al., 2024) and PyramidKV (Zhang et al., 2024) modify the KV cache capacity across different layers, prioritizing larger allocations in the lower layers while reducing those in the upper layers. However, most works in this category cannot be applied into training phase.
@@ 3.1 Fundamental Design
Separator Tokens. From Figure 2, we can observe that within a given input context, seemingly “meaningless” separator tokens (such as commas, periods, exclamation marks, semicolons, etc.) that segment sequences receive higher attention scores compared to semantically meaningful tokens (such as nouns or verbs). Therefore, we hypothesize that these separators may compress the information of the text segments naturally segmented by them, such that when the Transformer generates new tokens, it only needs to reference the information contained in these separators to extract the information pertaining to those text segments. Hence, in a training-free scenario, we employed this strategy and achieved similar results to the original model based on full attention across many tasks. Furthermore, to reinforce the effect of using separators to compress information within their respective segments, we employed training-from-scratch and post-training approaches to compel the model during training to restrict the current token from accessing all information from distant preceding text, i.e., in each segment, only the separator representing its segment is visible to the current token (with other tokens being masked, see Figure 3). After training in this manner, the information within segments is forced to be condensed into the separators, leading the Transformer’s probability distribution for predicting the next word closely resembling that of the original Transformer with full attention. See more in Appendices H and I.
Neighboring Tokens. Language tasks usually exhibit strong local dependencies and interactions, since adjacent tokens often form coherent phrases or have dependencies that are required to be captured. Neighboring tokens usually help form locally smooth and coherent contexts, allowing the model to generate sentences that are reasonable within the immediate context. The neighboring tokens, also referred to as local attention or sliding-window attention, are considered in various efficient Transformers (Xiao et al., 2024b; Zhang et al., 2023) and we have also adopted this approach, with the number of preceding tokens closest to the current token denoted as “n”.
@@ 3.2 Overall Pipeline
We split the overall pipeline of our proposed SepLLM into training/pre-filling stage and generating stage. We also provide a theoretical analysis about the Universal Approximation of SepLLM in Appendices J and K.
@@ 4.1 Experimental Settings
Table 1: Evaluation results and average runtime KV cache usage for training-free experiments on GSM8K-CoT 8-shots and MMLU 5-shots. For SepLLM and StreamingLLM, three initial tokens’ KV are kept for this experiment. r.KV (%) here represents the ratio of KV usage at runtime for the respective method compared to Vanilla. See more results in Appendices I and Table 17.
@@ 4.2 Training-Free
Method ARC-c ARC-e LBD-ppl LBD-acc LogiQA PIQA SciQ Atten. (%) r.KV (%)
@@ Results.
The experimental results for training-free are shown in Table 1. “Vanilla” represents the original Llama-3 model with full attention, while “StrmLLM” represents StreamingLLM (Xiao et al., 2024b)n means the number of KV for Neighboring Tokens we retain. For SepLLM, all the KV for Separator Tokens are kept and for the setting SepLLM (n=256), we find that SepLLM exhibits comparable performance in both multi-step mathematical CoT task and multidisciplinary knowledge reasoning tasks, when compared to the full-attention Llama-3. SepLLM achieves this using only 47.36% of the KV utilized by the original Llama-3 for reasoning, indicating SepLLM’s capability of modeling both the contexts requiring multi-step logical analysis and those involving multi-domain knowledge reasoning while retaining only 50% original KV.
@@ 4.6 Ablation Study
Settings.
We also perform the following experiments to validate the effectiveness of Initial Tokens and Positional Encoding’s shifting for both StreamingLLM and SepLLM. The experimental results are shown in Table 8 and the discussions are as follows.
@@
4.7 Comparison with More Baselines and Variants
Naive Baseline.
We compare SepLLM with more baselines and variants to verify its effectiveness. The results shown in Table 9 are based on a naive baseline in which $H_{s}$ out of $H$ attention heads are configured as sliding window attention, while the remaining heads use full attention. We use MMLU as the benchmark to measure the differences in reasoning ability, ensuring that the runtime KV usage for all methods is kept as consistent as possible. It can be observed that this naive head-wise baseline performs poorly (regardless of how many full attention heads are retained), which is due to the heterogeneity among the heads. In contrast, SepLLM achieves performance very close to Vanilla across various knowledge domains with the same KV usage.
State-of-the-Art Baselines.
In addition, we compare SepLLM with multiple state-of-the-art baseline methods, which demonstrates the effectiveness and simplicity of SepLLM. For detailed results, please refer to the Appendix E.
Fixed-Interval Variant.
Furthermore, to demonstrate the compression and summarization effects of separators on the information within the segments they divide, we propose a fixed-interval variant, FixLLM. In FixLLM, instead of attending to separator tokens, it attends to one token at fixed intervals, except for the Initial Tokens’ and Neighboring Tokens’ sections. We find that SepLLM significantly outperforms FixLLM, which indicates that the compression and summarization capabilities of separators in SepLLM cannot be replaced by fixed-interval tokens. Please see the Appendix I for detailed results and discussions.
@@ 4.8 Generalization and Information Retrieval
To verify the generalization of SepLLM, we adapt SepLLM to models of different architectures and scales. The results and discussions in Appendix D can validate the generalization capability of our proposed SepLLM. Specifically, we adapt SepLLM to different backbones including Pythia-6.9B, Pythia-12B (Biderman et al., 2023), Llama-3-8B-Base/Instruct (Dubey et al., 2024) and Falcon-40B (Almazrouei et al., 2023). Moreover, we also conduct the Needle-in-a-Haystack experiment, which further demonstrates the compression capability of separator tokens for segment information. As illustrated in Appendix F, SepLLM can retrieve the needle in most scenarios. In comparison, StreamingLLM (Xiao et al., 2024b) cannot complete this task. In addition, we provide comparisons of SepLLM with more baseline models regarding mathematical reasoning and logical analysis capabilities in the Appendix E. Furthermore, Appendices H and I discuss in detail the effects and functions of separators from the perspective of implementation logic.
MMLU
$H_{s}$/$H$ humanity social stem other overall r.KV (%) n
20/32 23.93 23.07 24.42 23.53 23.76 44.74 80
24/32 24.23 23.30 26.36 23.37 24.31 47.71 208
28/32 25.66 27.29 26.31 23.69 25.81 45.13 256
30/32 27.29 25.09 27.78 38.11 29.31 45.42 288
SepLLM 57.66 76.21 56.49 72.19 64.68 44.61 256
Vanilla 60.49 76.50 56.61 72.19 65.72 100.00 ALL
Table 9: Comparison with a naive baseline method, where $H_{s}$ out of $H$ attention heads are set to sliding window attention, and the remaining heads use full attention. n represents the number of Neighboring Tokens (i.e., the sliding window size). r.KV (%) indicates the ratio of KV usage at runtime for the respective method compared to Vanilla.
5 Universal Approximation
In this section, we present the theoretical results on the universal approximation capabilities of encoder-based SepLLM. Details can be found in Appendices J and K.
[…]
Theorem 5.1.
6 Native Sparse Attention and Language Modeling
SepLLM essentially models sparsity based on the natural semantic distribution of natural language. During the pre-training phase, SepLLM intentionally compresses segment information into the separator used to divide the segment. This approach closely aligns with the semantic distribution of natural language because the separator itself provides a division and summary of the current segment. The segments separated out are inherently semantically coherent, forming self-contained semantic units. Therefore, we can view SepLLM as a native sparse attention mechanism inherent to natural language itself. Moreover, SepLLM is suitable to replace StreamingLLM (which only statically focuses on important positions and is overly sparse) as the fundamental baseline model for sparse attention mechanisms in LLMs.
@@ 7 Concluding Remarks
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix D The Performance of Different Models
Appendix E Extended Comparisons
We use GSM8K-CoT (Cobbe et al., 2021), the most commonly used metric for testing mathematical reasoning and logical analysis, to compare SepLLM (a=3, n=256) with other state-of-the-art training-free methods, including H2(Zhang et al., 2023), SnapKV (Li et al., 2024), and PyramidKV (Zhang et al., 2024). All methods are configured to retain nearly identical runtime KV cache usage. The results are shown in Table 15. It can be observed that SepLLM, without requiring complex importance evaluation mechanisms, is able to maintain strong reasoning capabilities simply by compressing segment information.
flexible-extract strict-match r.KV (%)
Vanilla 77.79 77.26 100.00
SepLLM 77.18 77.18 47.36
H2O 76.27 75.06 47.54
SnapKV 76.50 73.62 47.54
PyramidKV 75.82 72.02 47.54
Table 15: Evaluation results and average runtime KV cache usage for experiments on GSM8K-CoT with 8-shots, compared to multiple baseline methods.
Appendix F Needle in a Haystack
@@
Figure 8: Needle-in-a-Haystack test results for StreamingLLM (n=64) based on Pythia-160M-deduped.
Figure 9: Needle-in-a-Haystack test results for our SepLLM(n=64, H/T) based on Pythia-160M-deduped.
Figure 10: Needle-in-a-Haystack test results for our SepLLM(n=2048; first/last 2 layers (4 layers in total): full attention) based on Llama-3-8B-instruct. 4 initial tokens are kept.
Figure 11: Needle-in-a-Haystack test results for our SepLLM(n=2048; first/last 2 layers (4 layers in total): full attention) based on Llama-3-8B-instruct. 32 initial tokens are kept.
@@
Appendix G Choice of Separators
In this paper, we treat the choice of separators as a hyperparameter. Since the number of commonly used separators is typically limited in most languages, we recommend including all types of separators as much as possible. In Table 16, we present the experimental results based on different choices of separators. We use the GSM8K-CoT metric to reflect the performance differences caused by these choices (GSM8K-CoT is a commonly used metric for evaluating the mathematical reasoning and logical analysis capabilities of LLMs). It can be observed that when using four common separators instead of the default nine, the performance shows a certain degree of decline. Furthermore, when only “.” and “?” are used, the model’s reasoning ability drops significantly. When all separators are removed, this performance degradation becomes even more pronounced (i.e., StrmLLM(n=256)). While increasing the number of Neighboring Tokens (n) to ensure the total amount of retained KV cache remains unchanged can partially mitigate the performance loss (i.e., StrmLLM(n=380)), it falls far short of restoring the performance to the levels achieved with the default nine separators or the four common separators.
Appendix H Discussions on Separators
@@
SepLLM (n=256) SepLLM (n=256) SepLLM (n=256) StrmLLM (n=256) StrmLLM (n=380)
[…]
r.KV (%) 47.36 37.92 36.44 31.92 47.54
flexible-extract 77.18 76.68 70.66 68.84 71.42
strict-match 77.18 76.85 70.36 67.63 70.89
Table 16: Evaluation results and average runtime KV cache usage for training-free experiments on GSM8K-CoT with 8-shots, based on different choices of separators.
Appendix I Fixed-Interval Variant
Sparsity is ubiquitous across various types of neural networks (Shi et al., 2021; Chen et al., 2025a; Zheng et al., 2025; Yuan et al., 2025; Chen et al., 2025b). To verify the sparsity caused by separators as natural semantic divisions in natural language and the resulting importance of these separators, we propose another variant, i.e., when calculating sparse attention, instead of focusing only on Separator Tokens between the Initial Tokens and Neighboring Tokens, we attend to one token at fixed intervals (e.g., every 8 tokens or 16 tokens), while the other tokens are masked, namely FixLLM. We evaluated the mathematical reasoning ability and knowledge-based reasoning ability of the two variants using the GSM8K-CoT (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021)) benchmarks under a training-free setting, based on the Llama3-8B-Instruct backbone. The results in Table 17 show that FixLLM has a significant gap compared to SepLLM in both mathematical logical reasoning and knowledge-based reasoning capabilities. Therefore, the summarization and compression effects of separators in SepLLM for the segments they divide cannot be replaced by fixed-interval tokens.
@@ Appendix I Fixed-Interval Variant
Table 17: Evaluation results and average runtime KV cache usage for training-free experiments on GSM8K-CoT 8-shots and MMLU 5-shots. For SepLLM and FixLLM, three initial tokens’ KV are kept. $\Delta$ denotes the interval size for FixLLM and n is the number of retained neighboring tokens’ KV. r.KV (%) represents the ratio of KV usage at runtime for the respective method compared to Vanilla.
Appendix J Universal Approximation
@@
Appendix K Proof for Theorem J.1
Lemma K.1 (Lemma 5 in Yun et al. (2020)).
@@
Definition K.3 (Contextual Mapping).
@@
Lemma K.5 (Lemma 8 in Yun et al. (2020)).
For the contextual mapping $g_{c}$ in Lemma K.4, there exist $\frac{n}{\delta^{dn}}$-layer modified SepLLM (composed of feed-forward layers and identity attention layers), denoted as $g_{v}\in\bar{\mathcal{T}}^{2,1,1}_{\text{Sep}}:\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times n}$, such that